OpenStudy (anonymous):

find the critical points for f(x)=cos^2(x) on [-pi,pi]

OpenStudy (anonymous):

hey r u there

OpenStudy (anonymous):

f'(x)=2(cos(x))(-sin(x)) for critical points set f'(x)=0 cos(x)=0 x=pi/2 and -sin(x)=0 x=0,pi/2

OpenStudy (anonymous):

x=pi/2,0,-pi/2